OpenStudy (anonymous):

A ball is thrown vertically upward from the roof of a building 60 ft high, so that its distance from the roof at any time t, in seconds, is given by the equation s=-16t^2+52t. Determine after how many seconds the ball will hit the roof.

OpenStudy (dumbcow):

set equal to 0...solve for t -16t^2 +52t = 0 factor out a t t(-16t+52) = 0 16t = 52 t = 52/16 = 13/4

OpenStudy (anonymous):

Thank you so much

OpenStudy (dumbcow):

yw